OpenStudy (anonymous):

Solve the inequality -3x + 4 x + 8 and show all of your work.

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

?

OpenStudy (anonymous):

missing something in the question

OpenStudy (anonymous):

Oh yeah. One second.

OpenStudy (anonymous):

Solve an inequality the same way you would solve a regular equation. With this ONE exception: If you multiply or divide both sides by a NEGATIVE number, then you should change the direction of the inequality sign.

OpenStudy (anonymous):

\(<, >, \geq , \leq\)

OpenStudy (anonymous):

easy peasy parchesie. =D

OpenStudy (anonymous):

Solve the inequality -3x + 4<= x + 8 and show all of your work.

OpenStudy (anonymous):

fortunately you never have to divide by a negative number

OpenStudy (anonymous):

Give it a try on your own and ask me any questions that come up =)

OpenStudy (anonymous):

What's the first step? lol

OpenStudy (anonymous):

Group like terms. Get x by itself.

OpenStudy (anonymous):

Don't be intimidated. It's nearly the same as solving a normal equation.

OpenStudy (anonymous):

add 4 and 8?

OpenStudy (anonymous):

\[-3x+4\leq x+8\] add \(3x\) to both sides and get \[4\leq 4x+8\]

OpenStudy (anonymous):

Get them on the same side first.

OpenStudy (anonymous):

Now do I divide 4x by 4?

OpenStudy (anonymous):

subtract 4 from 8 on bothsides then take 1 away from the negative 3x to get -2 then divide 4 by negative two to get...\[x \le -2\]

OpenStudy (anonymous):

can you write all that out?

OpenStudy (anonymous):

When your write it in words it confuses me.

OpenStudy (anonymous):

yeah \[-3x+4 \le x+8\\[x \le -2\] \[-3x \le x + 4\] \[-2x \le 4\

OpenStudy (anonymous):

−3x+4 ≤ x+8 First of all, these steps are exactly the same as if we were solving the problem −3x+4 = x+8 The first step is to get like terms together. That means I want x on one side and everything else on the other. −3x+4 ≤ x+8 Add 3x to both sides. 4 ≤ 4x+8 Subtract 8 from both sides. -4 ≤ 4x divide both sides by 4.

OpenStudy (anonymous):

That is explains it a lot better. thank you(:

OpenStudy (anonymous):

I think you're letting yourself get scared, but solving this kind of problem is almost exactly the same as normal equations. There is one little difference, and I want to show it to you.

OpenStudy (anonymous):

-1<=x is the answer?

OpenStudy (anonymous):

Yes, that's right. If you ever multiply or divide both sides by a negative number, you have to switch which way the inequality points. I looks like this: -2x < 6 divide both sides by (-2), and remember to switch the inequality sign x > -3

OpenStudy (anonymous):

Can you help me with another one?. please.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve the compound inequality 6x + 7 > 1 and 6x + 7 < 25 and show all of your work.

OpenStudy (anonymous):

Solve each of them separately. Please show me your attempt.

OpenStudy (anonymous):

You can do this. Trust me.

OpenStudy (anonymous):

Divide 6 from 7?

OpenStudy (anonymous):

Not yet. Get like terms together first.

OpenStudy (anonymous):

7 minus 1?

OpenStudy (anonymous):

Start with the actual whole inequality, do something to both sides, and show me the result. You're just making guesses -_-

OpenStudy (anonymous):

No I'm not You said to start with the separately.. I don't get this. I honestly.

OpenStudy (anonymous):

Please help me.

OpenStudy (anonymous):

It's alright. Let's try some simple equations. I'll do an example, and then I'll give you one to try. Example: 3x + 1 = 7 I start by getting like terms together. I have 1 on the left, and 7 on the right, so I need to move 1 over to the left. How do I do that? Subtract 1 from both sides. What is the result when I do that? 3x + 1 - 1 = 7-1 which gives 3x = 6 Now, I have like terms grouped, but x is not by itself yet. x is multiplied by 3, so how do I get rid of it? Divide both sides by 3, which gives me 3x/3 = 6/3 x = 2

OpenStudy (anonymous):

Now you try this one. 5x - 6 = 9

OpenStudy (anonymous):

I need help with the problem I have given.

OpenStudy (anonymous):

Right, but you don't know how to solve basic equations, so the best way for me to help you is to teach you that.

OpenStudy (anonymous):

Are you here to learn, or are you here to get answers?

OpenStudy (anonymous):

Is this right? Part 1: 6x +7>1 6x>-6 x>-1 6x + 7 < 25 6x<18 x<3 -1<x<3

OpenStudy (anonymous):

Yes =)

OpenStudy (anonymous):

And that is exactly what I'm looking for. Change the inequality one step at a time until it is solved for x.

OpenStudy (anonymous):

How would graph that on an number line?

OpenStudy (anonymous):

CAN YOU People tell me whats exactly the question?

OpenStudy (anonymous):

Would it be a open circle on 1 then colored to the right connected to a open circle on 3 colored to the left?

OpenStudy (anonymous):

Good =) Except be careful. Not 1, but -1. Right?

OpenStudy (anonymous):

Oh yeah. I meant that.

OpenStudy (anonymous):